OpenStudy (anonymous):

I think THIS is the chain rule Let y = u**(9/2) and u = 8 x + 1. what is the derivitive?

OpenStudy (anonymous):

** means ^

OpenStudy (anonymous):

dy/dx = dy/du * du/dx

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i mean dy/dx=8

OpenStudy (anonymous):

dy/du = (9/2) u ^ (7/2) now plug in u = 8x + 1 and multiply by 8 du/dx = 8

OpenStudy (anonymous):

dy/du=9/2x^1-9/2

OpenStudy (anonymous):

wait do i use the chain rule?

OpenStudy (anonymous):

64x+8

OpenStudy (anonymous):

i think 9/2u^7/2 is wrong

OpenStudy (anonymous):

if y = u^(9/2) dy/du = (9/2 ) u^ (9/2 - 1 ) this is the power rule: if y=ax^n dy/du = anx^(n-1)

OpenStudy (anonymous):

i did that but i think 9/2u^7/2 is wrong

OpenStudy (anonymous):

you do use chain rule dy/dx = dy/du * du/dx

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i thought you do the power rule inside the chain rule since its in brackets

OpenStudy (anonymous):

ok y = u^(9/2) and u = 8x + 1 right?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

dy/dx = (9/2) u ^ (7/2) * 8 = (9/2) (8x + 1)^(7/2) * 8 ( plugging in u = 8x+1) = 36 ( 8x + 1)^(7/2)

OpenStudy (anonymous):

theres something wrong w/ dy/du

OpenStudy (anonymous):

thats the only problem because du/dx is right

OpenStudy (anonymous):

so the process is f=right but one number is wrong

OpenStudy (anonymous):

lets look at y = x^n them dy/dx = n x ^ (n-1) now let u = x and n = 9/2 then y = u^(9/2) so dy/du = (9/2) u ^(9/2 - 1) = (9/2) u^ (7/2) because 9/2 - 1 = 9/2 - 2/2 = 7/2